Be sure to answer all parts. The pH of an aqueous acid solution is 6.24 at 25°C. Calculate the Ka for the acid. The initial acid concentration is 0.010 M. × 10 (Enter your answer in scientific notation.)
Answer #1

HA solation: HA + 40 40 + A- kan (tegex] GA-7 o given - Hal=0.000m pre 6-24 , Hence, (rizo+7= 10ph or sotzet) = 10 (6-24) = 5